### 13 Mar 2026 v1

# Three elliptic closed characteristics on the non-degenerate compact convex hypersurfaces in $\mathbb{R}^{6}$

Abstract
Let $\Sigma\subset\mathbb{R}^{2n}$ with $n\geq 2$ be any $C^{2}$ compact convex hypersurface. The stability of closed characteristics has attracted considerable attention in related research fields. A long-standing conjecture states that all closed characteristics are irrationally elliptic, provided $\Sigma$ possesses only finitely geometrically distinct closed characteristics. This conjecture has been fully resolved only in $\mathbb{R}^{4}$, while it remains completely open in higher dimensions. Even in $\mathbb{R}^{6}$, it is unknown whether there exist three elliptic closed characteristics. In this paper, we first prove that for any non-degenerate $\Sigma\in\mathbb{R}^{2n}$ with finitely many closed characteristics, at least two of them are irrationally elliptic. Moreover, for any non-degenerate $\Sigma\subset\mathbb{R}^{6}$ with finitely many closed characteristics, we obtain at least three elliptic characteristics, of which at least two are irrationally elliptic. Based on the $n$-or-$\infty$ conjecture, three elliptic closed characteristics are optimal. This result provide theoretical support for further research on this conjecture.
AMS Subject Classification: 58E05, 37J45, 34C25
Key Words: Compact convex hypersurfaces, Closed characteristics, Stability, Maslov-type index, Hamiltonian systems
1 Introduction and main results
Let $\Sigma$ be a $C^{2}$ compact hypersurface in $\mathbb{R}^{2n}$ that is strictly convex with respect to the origin, i.e., the tangent hyperplane of $\Sigma$ at any $x\in\Sigma$ does not pass through the origin. Without loss of generality, we suppose it contains the origin. We denote the set of all such hypersurfaces in $\mathbb{R}^{2n}$ by $\mathcal{H}(2n)$.
For $x\in\Sigma$, let $N_{\Sigma}(x)$ be the outward unit normal vector of $\Sigma$ at $x$, satisfying $N_{\Sigma}(x)\cdot x=1$. We consider closed characteristics $(\tau,x)$ on $\Sigma$, which are solutions of the following problem
$\begin{cases}\dot{x}(t)=JN_{\Sigma}(x(t)),\ x(t)\in\Sigma,\ \forall t\in\mathbb{R},\\ x(\tau)=x(0),\end{cases}$ (1.1)
where $J=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}$ is the standard symplectic matrix and $\tau>0$. A closed characteristic $(\tau,x)$ is prime with $\tau$ being the minimal period. Two closed characteristics $(\tau,x)$ and $(\sigma,y)$ are called geometrically distinct if $x(\mathbb{R})\neq y(\mathbb{R})$. The set of all geometrically distinct closed characteristics is denoted by $\widetilde{\mathcal{J}}(\Sigma)$, and let $(\tau,x)]$ be the equivalence class of closed characteristics geometrically identical to $(\tau,x)$.
Let $j:\mathbb{R}^{2n}\to\mathbb{R}$ be the gauge function of $\Sigma$, i.e., $j(\lambda x)=\lambda$ for $x\in\Sigma$ and $\lambda\geq 0$, then $j\in C^{2}(\mathbb{R}^{2n}\setminus\{0\},\mathbb{R})\cap C^{1}(\mathbb{R}^{2n},\mathbb{R})$ and $\Sigma=j^{-1}(1)$. Fix a constant $\alpha\in(1,2)$ and define the Hamiltonian function $H_{\alpha}:\mathbb{R}^{2n}\to[0,+\infty]$ by
$H_{\alpha}(x)=j(x)^{\alpha},\quad\forall x\in\mathbb{R}^{2n}.$
Then $H_{\alpha}\in C^{2}(\mathbb{R}^{2n}\setminus\{0\},\mathbb{R})\cap C^{1}(\mathbb{R}^{2n},\mathbb{R})$ is convex and $\Sigma=H_{\alpha}^{-1}(1)$. It is well known that the problem (1.1) is equivalent to the following given energy problem of the Hamiltonian system
$\begin{cases}\dot{x}(t)=JH_{\alpha}^{\prime}(x(t)),\ H_{\alpha}(x(t))=1,&\forall t\in\mathbb{R},\\ x(\tau)=x(0).\end{cases}$ (1.2)
Denote by $\widetilde{\mathcal{J}}(\Sigma,\alpha)$ the set of all geometrically distinct solutions $(\tau,x)$ of (1.2) where $\tau$ is the minimal period of $x$. Note that elements in $\widetilde{\mathcal{J}}(\Sigma)$ and $\widetilde{\mathcal{J}}(\Sigma,\alpha)$ are one to one correspondent to each other.
Let $(\tau,x)\in\widetilde{\mathcal{J}}(\Sigma,\alpha)$, the fundamental solution $\gamma_{x}:[0,\tau]\to\text{Sp}(2n)$ with $\gamma_{x}(0)=I_{2n}$ of the linearized Hamiltonian system
$\dot{y}(t)=JH_{\alpha}^{\prime\prime}(x(t))y(t),\quad\forall t\in\mathbb{R}$ (1.3)
is called the associated symplectic path of $(\tau,x)$. The eigenvalues of $\gamma_{x}(\tau)$ are called Floquet multipliers, and denote the elliptic height $e(\gamma_{x}(\tau))$ of $\gamma_{x}(\tau)$ to be the total algebraic multiplicity of all eigenvalues of $\gamma_{x}(\tau)$ on the unit circle $\mathbb{U}=\{z\in\mathbb{C}\,|\,|z|=1\}$. As usual, a closed characteristic $(\tau,x)$ is non-degenerate, if $1\in\sigma(x)$ is of algebraic multiplicity $2$; a hypersurface $\Sigma$ is called non-degenerate, if all the closed characteristics on the given hypersurface together with all of their iterations are non-degenerate as periodic solutions of the corresponding Hamiltonian system. A closed characteristic $(\tau,x)$ is called elliptic, if $e(\gamma_{x}(\tau))=2n$; hyperbolic if $1$ is a double Floquet multiplier of it and $e(\gamma_{x}(\tau))=2$; irrationally elliptic if in an appropriate coordinate $\gamma_{x}(\tau)$ is the $\diamond$-product of one $N_{1}(1,1)=\left(\begin{array}[]{cc}1&1\\ 0&1\\ \end{array}\right)$ and n1n-1 rotation matrices $\left(\begin{array}[]{cc}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\\ \end{array}\right)$ with rotation angles $\theta_{i}$ being irrational multiples of $\pi$ for $1\leq i\leq n-1$. Hence, the stability problem refers to estimating the number of elliptic closed characteristics that may exist for system (1.3).
Research on the stability of closed characteristics dates back to 1986, I. Ekeland [6] proved that if $\Sigma\in\mathcal{H}(2n)$ is a $\sqrt{2}$-pinched compact convex hypersurface, then it admits at least one elliptic closed characteristic. Later, he further made following conjectured in [7]:
Conjecture 1. For any $\Sigma\in\mathcal{H}(2n)$, there exists at least one elliptic closed characteristic.
In [5], Dell’Antonio-D’Onofrio-Ekeland verified the above conjecture by imposing the symmetry condition $\Sigma=-\Sigma$. In 1998, Long [15] showed that for any $\Sigma\in\mathcal{H}(2n)$, either there exist infinitely many geometrically distinct hyperbolic closed characteristics $(\tau_{j},x_{j})]$ with minimal periods $\tau_{j}\to+\infty$, or there exists at least one non-hyperbolic closed characteristic. In 2002, Long and Zhu [19] refined this result: If $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then there has at least one elliptic closed characteristic on $\Sigma$, and at least $\frac{n}{2}]$ geometrically distinct closed characteristics possess irrational mean indices, where #A denotes the total number of elements in a set A. They further proved that if $^{\#}\widetilde{\mathcal{J}}(\Sigma)\leq 2\rho_{n}(\Sigma)-2<+\infty$, then $\Sigma$ admits at least two elliptic closed characteristics. Also, in 2000, Long [16] proved that for $n=2$, if $^{\#}\widetilde{\mathcal{J}}(\Sigma)=2$, then both closed characteristics are elliptic. Later, in 2007, [22] strengthened this conclusion by showing that they are in fact irrational elliptic closed characteristics. Based on these results, when $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, it is natural to guess that we should have a stronger result than the conclusion of Conjecture 1. Therefore, in [22], the authors further made the following conjecture:
Conjecture 2. For any $\Sigma\in\mathcal{H}(2n)$, assume $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then all the closed characteristics on $\Sigma$ are irrationally elliptic.
A typical example is the non-resonant ellipsoid in $\mathbb{R}^{2n}$, that is $\Sigma$ is defined by
$\sum_{i=1}^{n}\frac{\alpha_{i}}{2}(p_{i}^{2}+q_{i}^{2})=1,$
where $\alpha_{i}/\alpha_{j}\in\mathbb{R}\setminus\mathbb{Q}$. There just exist $n$ closed characteristics $x_{i},i=1,...,n$ and their mean Maslov-type index satisfy $\hat{i}(x_{i})/\hat{i}(x_{j})=\alpha_{j}/\alpha_{i}\in\mathbb{R}\setminus\mathbb{Q}$ and all the closed characteristics are irrationally elliptic. Regarding the progress on Conjecture 2, in 2006, Long-Wang [18] proved that for $n=3$, if $^{\#}\widetilde{\mathcal{J}}(\Sigma)=3$, then at least two of the closed characteristics are elliptic. In 2014, Wang [20] further upgraded this result to show that there exist at least two irrational elliptic closed characteristics. In 2022, Wang [21] established that if $\rho_{n}(\Sigma)=n$ and $^{\#}\widetilde{\mathcal{J}}(\Sigma)=n$, then at least two of them are irrational elliptic closed characteristics. Recently, Li-Liu-Wang [11] improves [21] by eliminating the technical condition $\rho_{n}(\Sigma)=n$ and prove that if $^{\#}\widetilde{\mathcal{J}}(\Sigma)=n$, then at least two of them are irrational elliptic closed characteristics. For results on more general compact star-shaped hypersurfaces, we refer the reader to [12], [9], [13], [11], etc.
Remark 1.1.
In general, when $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, it hard to estimate the upper bound of $^{\#}\widetilde{\mathcal{J}}(\Sigma)$. This relate to the following conjecture.
The $n$-or-$\infty$ conjecture: For any $\Sigma\in\mathcal{H}(2n)$, then $^{\#}\widetilde{\mathcal{J}}(\Sigma)=n$ or $\infty$.
In [8] of 1998, Hofer-Wysocki-Zehnder proved that $^{\#}\widetilde{\mathcal{J}}(\Sigma)=2$ or $\infty$ holds for every $\Sigma\in\mathcal{H}(4)$. Furthermore, Colin-Dehornoy-Rechtman [3] proved that there are two or infinitely many simple Reeb orbits for any non-degenerate contact form on a closed connected three-manifold. Recently, this conjecture was confirmed in [4], by proving a stronger theorem that the two or infinity result holds for any closed contact 3-manifold whose contact structure possessing torsion first Chern class. However, for $n\geq 3$, this conjecture is still widely open. It is worth mentioning that recent progress by Çineli-Ginzburg-Gürel in [1] of 2024 confirmed that $^{\#}\widetilde{\mathcal{J}}(\Sigma)\geq n$ for $\Sigma$ bounding star-shaped domains with dynamically convex Reeb flows.
Since the upper bound of $^{\#}\widetilde{\mathcal{J}}(\Sigma)$ is hard to estimate, as it involves the challenging the $n$-or-$\infty$ conjecture. In 2017, under the original finiteness condition in Conjecture 2, i.e $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, Hu-Ou [10] proved that for any $\Sigma\in\mathcal{H}(2n)$, there has at least two elliptic closed characteristics on $\Sigma$. In this paper, we further obtain the following conclusions.
Theorem 1.2.
For any non-degenerate surfaces $\Sigma\in\mathcal{H}(2n)$ with $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then at least two closed characteristics are irrational elliptic.
Remark 1.3.
By combining the methods of [1], [11] and this paper, the result of Theorem 1.2 can perhaps be generalized to dynamically convex Reeb flow on non-degenerate star-shaped surfaces.
Remark 1.4.
How to remove the non-degeneracy condition in Theorem 1.2 is an interesting question, but for now, we are unable to do so. In fact, it is still unknown whether the following Assumption A holds, except for the case $n=2$. We believe it to be true for $n\geq 3$, but we are currently unable to prove it; perhaps a deeper analysis of the methods in [11] might lead to progress.
Assumption A: Consider the surfaces $\Sigma\in\mathcal{H}(2n)$ with $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then for any $a\in A(v)$, $x_{j(1)}$ is an irrational elliptic closed characteristics, where $A(v)$ and $x_{j(1)}$ is given by Theorem 2.4.
Theorem 1.5.
Under Assumption A, then for any surfaces $\Sigma\in\mathcal{H}(2n)$ with $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, at least two closed characteristics are irrational elliptic.
So far, Conjecture 2 has been fully resolved only in $\mathbb{R}^{4}$, while it remains entirely an open problem in higher dimensions. Even in $\mathbb{R}^{6}$, it is unknown whether there exist three elliptic closed characteristics. Therefore, we further prove that,
Theorem 1.6.
For any non-degenerate surfaces $\Sigma\in\mathcal{H}(6)$ with $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then at least three closed characteristics are elliptic, and among these three, at least two are irrational elliptic.
Remark 1.7.
Based on the $n$-or-$\infty$ conjecture, when $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then $^{\#}\widetilde{\mathcal{J}}(\Sigma)=3$, hence three elliptic closed characteristics is optimal.
Our results are mainly derived through analytical deduction by using of the Long-Zhu’s common index jump methods [19] and the estimation methods on iterative indices presented in the work of Hu-Ou [10]. In Section 2, we introduce the relevant concepts and formulas involved in this paper, and Section 3 provides a concise proof of the main Theorems.
[…]
$\begin{cases}{\lfloor a\rfloor}=\max\{k\in\mathbb{Z}\mid k\leq a\},\quad&E(a)=\min\{k\in\mathbb{Z}\mid k\geq a\},\\ \{a\}=a-\lfloor a\rfloor,\quad&\varphi(a)=E(a)-\lfloor a\rfloor.\end{cases}$
2 Review of the Maslov-type index theory
In this section, we briefly review the index theory for symplectic paths, for comprehensive details, please refer to [14], [16] and [19].
2.1 Definition of the Maslov-type index
As usual, let $\omega\in\mathbb{U}$, the symplectic group $\mathrm{Sp}(2n)$ is defined by
$\mathrm{Sp}(2n)=\{M\in\mathrm{GL}(2n,\mathbb{R})\,|\,M^{T}JM=J\},$
whose topology is induced from that of $\mathbb{R}^{4n^{2}}$. For $\tau>0$ we are interested in paths in $\mathrm{Sp}(2n)$:
$\mathcal{P}_{\tau}(2n)=\{\gamma\in C([0,\tau],\mathrm{Sp}(2n))\,|\,\gamma(0)=I_{2n}\},$
which is equipped with the topology induced from that of $\mathrm{Sp}(2n)$. The following real function was introduced in [16]:
$D_{\omega}(M)=(-1)^{n-1}\overline{\omega}^{n}\det(M-\omega I_{2n}),\ \ \forall\omega\in\mathbb{U},\,M\in\mathrm{Sp}(2n).$
Thus for any $\omega\in\mathbb{U}$ the following codimension one hypersurface in $\mathrm{Sp}(2n)$ is defined in [16]:
$\mathrm{Sp}(2n)_{\omega}^{0}=\{M\in\mathrm{Sp}(2n)\,|\,D_{\omega}(M)=0\}.$
For any $M\in\mathrm{Sp}(2n)_{\omega}^{0}$, we define a co-orientation of $\mathrm{Sp}(2n)_{\omega}^{0}$ at $M$ by the positive direction $\frac{d}{dt}Me^{t\epsilon J}|_{t=0}$ of the path $Me^{t\epsilon J}$ with $0\leq t\leq 1$ and $\epsilon>0$ being sufficiently small. Let
$\mathrm{Sp}(2n)_{\omega}^{\ast}$ $=$ $\mathrm{Sp}(2n)\setminus\mathrm{Sp}(2n)_{\omega}^{0},$
$\mathcal{P}_{\tau,\omega}^{\ast}(2n)$ $=$ $\{\gamma\in\mathcal{P}_{\tau}(2n)\,|\,\gamma(\tau)\in\mathrm{Sp}(2n)_{\omega}^{\ast}\},$
$\mathcal{P}_{\tau,\omega}^{0}(2n)$ $=$ $\mathcal{P}_{\tau}(2n)\setminus\mathcal{P}_{\tau,\omega}^{\ast}(2n).$
For any two continuous arcs $\xi$ and $\eta:[0,\tau]\to\mathrm{Sp}(2n)$ with $\xi(\tau)=\eta(0)$, it is defined as usual:
$\eta\ast\xi(t)=\begin{cases}\xi(2t),&\text{if }0\leq t\leq\tau/2,\\ \eta(2t-\tau),&\text{if }\tau/2\leq t\leq\tau.\end{cases}\\$
Given any two $2m_{k}\times 2m_{k}$ matrices of square block form $M_{k}=\left(\begin{array}[]{cc}A_{k}&B_{k}\\ C_{k}&D_{k}\\ \end{array}\right)$ with $k=1,2$, as in [17], the $\diamond$-product of $M_{1}$ and $M_{2}$ is defined by the following $2(m_{1}+m_{2})\times 2(m_{1}+m_{2})$ matrix $M_{1}\diamond M_{2}$:
$M_{1}\diamond M_{2}=\left(\begin{array}[]{cccc}A_{1}&0&B_{1}&0\\ 0&A_{2}&0&B_{2}\\ C_{1}&0&D_{1}&0\\ 0&C_{2}&0&D_{2}\\ \end{array}\right),$
and denote by $M^{\diamond k}$ the $k$-fold $\diamond$-product $M\diamond\cdots\diamond M$. Note that the $\diamond$-product of any two symplectic matrices is symplectic. For any two paths $\gamma_{j}\in\mathcal{P}_{\tau}(2n_{j})$ with $j=0$ and $1$, let $\gamma_{0}\diamond\gamma_{1}(t)=\gamma_{0}(t)\diamond\gamma_{1}(t)$ for all $t\in[0,\tau]$.
For $\gamma(\tau)\in\mathrm{Sp}(2n)_{\omega}^{*}$, the $\omega$-index of $\gamma$ can be defined by
$i_{\omega}(\gamma)=\left[\mathrm{Sp}(2n)_{\omega}^{0}:\gamma*\xi_{n}\right],$ (2.1)
where $\xi_{n}(t)=\begin{pmatrix}2-\frac{t}{\tau}&0\\ 0&\left(2-\frac{t}{\tau}\right)^{-1}\end{pmatrix}^{\diamond n},0\leq t\leq\tau$ and the right hand side of (2.1) is the usual homotopy intersection number, the orientation of $\gamma\ast\xi_{n}$ is its positive time direction under homotopy with fixed end points.
For $\gamma(\tau)\in\mathrm{Sp}(2n)_{\omega}^{0}$, the $\omega$-index of $\gamma$ is defined by
$i_{\omega}(\gamma)=\sup_{U\in\mathcal{F}(\gamma)}\inf\left\{i_{\omega}(\beta)\mid\beta\in U\cap\mathcal{P}_{\tau,\omega}^{*}(2n)\right\},$
where $\mathcal{F}(\gamma)$ denotes the set of open neighborhoods of $\gamma$ in $\mathcal{P}_{\tau}(2n)$. For any symplectic matrix $M\in\mathrm{Sp}(2n)$, the $\omega$-nullity is
$\nu_{\omega}(M)=\dim_{\mathbb{C}}\ker_{\mathbb{C}}(M-\omega I).$
In particular, we set $\nu_{\omega}(\gamma):=\nu_{\omega}(\gamma(\tau))$. The integer pair $(i_{\omega}(\gamma),\nu_{\omega}(\gamma))\in\mathbb{Z}\times\{0,1,\dots,2n\}$ is called the index function of $\gamma$ with respect to $\omega$.
For any $\gamma\in\mathcal{P}_{\tau}(2n)$, For simplicity, define
$(i(\gamma,k),\nu(\gamma,k))=(i_{1}(\gamma^{k}),\nu_{1}(\gamma^{k})),\ \ \forall k\in\mathbb{N}.$
where $(i_{1}(\gamma^{k}),\nu_{1}(\gamma^{k}))$ is the index function of the iterative symplectic path $\gamma^{k}(t):=\gamma(t-j\tau)\gamma(\tau)^{j}$, here $j\tau\leq t\leq(j+1)\tau$, $j=0,1,\dots,k-1$. The mean index $\hat{i}(\gamma,m)$ per $m\tau$ for $m\in\mathbb{N}$ is defined by
$\hat{i}(\gamma,m)=\lim_{k\to+\infty}\frac{i(\gamma,mk)}{k}.$
For $\Sigma\in\mathcal{H}(2n)$ and $\alpha\in(1,2)$, let $(\tau,x)\in\mathcal{J}(\Sigma,\alpha)$. For simplicity, we define
$S^{+}(x)$ $=$ $S^{+}_{\gamma_{x}(\tau)}(1),$
$(i(x,m),\nu(x,m))$ $=$ $(i(\gamma_{x},m),\nu(\gamma_{x},m)),$
$\hat{i}(x,m)$ $=$ $\hat{i}(\gamma_{x},m),$
For all $m\in\mathbb{N}$, where $\gamma_{x}$ is the associated symplectic path of $(\tau,x)$.
2.2 Iterative formula of the Maslov-type index
The splitting numbers of a symplectic matrix $M\in\mathrm{Sp}(2n)$ with respect to $\omega\in\mathbb{U}$ can be defined as
$S_{M}^{\pm}(\omega)=\lim_{\varepsilon\to 0^{+}}i_{\omega\exp(\pm\sqrt{-1}\varepsilon)}(\gamma)-i_{\omega}(\gamma).$
Clearly, if $\omega\notin\sigma(M)$, then $S_{M}^{\pm}(\omega)=0$. If $\omega\in\sigma(M)$, since splitting numbers are symplectically additive and any symplectic matrix can be decomposed into the diamond product of basic normal forms, we present the splitting numbers of basic normal forms below.
For the basic normal form
$D(\lambda)=\begin{pmatrix}\lambda&0\\ 0&\lambda^{-1}\end{pmatrix},\quad\lambda=\pm 2;\ \ N_{1}(\lambda,a)=\begin{pmatrix}\lambda&a\\ 0&\lambda\end{pmatrix},\quad\lambda=\pm 1,a=\pm 1,0,$
$R(\theta)=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix},\ \ N_{2}(\omega,b)=\begin{pmatrix}R(\theta)&b\\ 0&R(\theta)\end{pmatrix},\quad\theta\in(0,\pi)\cup(\pi,2\pi),$
where $b=\begin{pmatrix}b_{1}&b_{2}\\ b_{3}&b_{4}\end{pmatrix}$, $b_{i}\in\mathbb{R}$ and $b_{2}\neq b_{3}$, $\omega=e^{\sqrt{-1}\theta}$, we have
$\left(S_{D(\lambda)}^{+}(e^{\sqrt{-1}\theta}),S_{D(\lambda)}^{-}(e^{\sqrt{-1}\theta})\right)$ $=(0,0),\ \ \left(S_{R(\theta)}^{+}(e^{\sqrt{-1}\theta}),S_{R(\theta)}^{-}(e^{\sqrt{-1}\theta})\right)=(0,1),$
[…]
Here we call $N_{2}(\omega,b)$ nontrivial if $(b_{2}-b_{3})\sin\theta<0$ and trivial if $(b_{2}-b_{3})\sin\theta>0$. For any $M_{i}\in\emph{Sp}(2n_{i})$ with $i=0$ and $1$, there holds
$S^{\pm}_{M_{0}\diamond M_{1}}(\omega)=S^{\pm}_{M_{0}}(\omega)+S^{\pm}_{M_{1}}(\omega),\ \ \text{and}\ \ S^{\pm}_{M}(\omega)=S^{\mp}_{M}(\bar{\omega}),\ \ \forall\ \ \omega\in\mathbb{U}.$
where $\bar{\omega}$ is the conjugate of $\omega$. Based on the splitting numbers, Long-Zhu [19] presented the abstract precise iteration formula for the Maslov-type index:
$i(\gamma,m)$ $=m\left(i(\gamma,1)+S_{M}^{+}(1)-C(M)\right)$
$\quad+2\sum_{\theta\in(0,2\pi)}E\left(\frac{m\theta}{2\pi}\right)S_{M}^{-}\left(e^{\sqrt{-1}\theta}\right)-\left(S_{M}^{+}(1)+C(M)\right),$
where $C(M)=\sum_{0<\theta<2\pi}S_{M}^{-}\left(e^{\sqrt{-1}\theta}\right),$ and $\gamma(\tau)$ is symplectically similar to the following form,
$\gamma(\tau)$ $\simeq N_{1}(1,1)^{\diamond p_{-}}\diamond I_{2p_{0}}\diamond N_{1}(1,-1)^{\diamond p_{+}}\diamond N_{1}(-1,1)^{\diamond q_{-}}\diamond-I_{2q_{0}}\diamond N_{1}(-1,-1)^{\diamond q_{+}}$ (2.2)
$\quad\diamond R(\theta_{1})\diamond\cdots\diamond R(\theta_{r})\diamond N_{2}(\omega_{1},u_{1})\diamond\cdots\diamond N_{2}(\omega_{r_{*}},u_{r_{*}})$
$\quad\diamond N_{2}(\lambda_{1},v_{1})\diamond\cdots\diamond N_{2}(\lambda_{r_{0}},v_{r_{0}})\diamond M_{k},$
where $N_{2}(\omega_{j},u_{j})$ are nontrivial forms with $\omega_{j}=e^{\sqrt{-1}\alpha_{j}}$, $\alpha_{j}\in(0,\pi)\cup(\pi,2\pi)$ and $u_{j}=\begin{pmatrix}u_{j1}&u_{j2}\\ u_{j3}&u_{j4}\end{pmatrix}\in\mathbb{R}^{2\times 2}$; $N_{2}(\lambda_{j},v_{j})$ are trivial forms with $\lambda_{j}=e^{\sqrt{-1}\beta_{j}}$, $\beta_{j}\in(0,\pi)\cup(\pi,2\pi)$ and $v_{j}=\begin{pmatrix}\nu_{j1}&\nu_{j2}\\ \nu_{j3}&\nu_{j4}\end{pmatrix}\in\mathbb{R}^{2\times 2}$; $M_{k}=D(2)^{\diamond k}$ or $D(-2)\diamond D(2)^{\diamond(k-1)}$; $p_{-},p_{0},p_{+},q_{-},q_{0},q_{+},r,r_{*}$ and $r_{0}$ are non-negative integers. Furthermore, we say that a normal form is a rational normal form if the rotation angle in the normal form is a rational multiple of $\pi$. Then the number of the rational normal form in $\{R(\theta_{1}),\ldots,R(\theta_{r})\}$ is denoted by $\tilde{r}$. Similarly, for set $\{N_{2}(\omega_{1},u_{1}),\ldots,N_{2}(\omega_{r_{*}},u_{r_{*}})\}$ and $\{N_{2}(\lambda_{1},\nu_{1}),\ldots,N_{2}(\lambda_{r_{0}},\nu_{r_{0}})\}$, the number of rational normal form is denoted by $\tilde{r}_{*}$ and $\tilde{r}_{0}$ respectively. These integers and real numbers are uniquely determined by $\gamma(\tau)$.
Then we can rewrite the iteration formula as
$i(\gamma,m)$ $=m\left(i(\gamma,1)+p_{-}+p_{0}-r\right)+2\sum_{l=1}^{r}E\left(\frac{m\theta_{j}}{2\pi}\right)-r-p_{-}-p_{0}$
$\quad-\frac{1+(-1)^{m}}{2}(q_{0}+q_{+})+2\left(\sum_{l=1}^{r_{*}}\varphi\left(\frac{m\alpha_{j}}{2\pi}\right)-r_{*}\right),$
$\nu(\gamma,m)$ $=\nu(\gamma,1)+\frac{1+(-1)^{m}}{2}(q_{-}+2q_{0}+q_{+})+2(r+r_{*}+r_{0})$ (2.3)
$\quad-2\left(\sum_{l=1}^{r}\varphi\left(\frac{m\theta_{l}}{2\pi}\right)+\sum_{l=1}^{r_{*}}\varphi\left(\frac{m\alpha_{l}}{2\pi}\right)+\sum_{l=1}^{r_{0}}\varphi\left(\frac{m\beta_{l}}{2\pi}\right)\right),$
$\hat{i}(\gamma,1)=i(\gamma,1)+p_{-}+p_{0}-r+\sum_{l=1}^{r}\frac{\theta_{l}}{\pi},$ (2.4)
$S_{M}^{+}(1)=p_{-}+p_{0},$ (2.5)
$C(M)=\sum_{0<\theta<2\pi}S_{M}^{-}(e^{\sqrt{-1}\theta})=q_{0}+q_{+}+r+2r_{*}.$ (2.6)
2.3 Common index jump theorem
We first introduce the common index jump theorem for finite families of symplectic paths, which can be found in Long-Zhu [19].
Theorem 2.1.
Let $\gamma_{k}\in\mathcal{P}_{\tau_{k}}(2n)$ for $k=1,\dots,q$ be a finite family of symplectic paths, and let $M_{k}=\gamma_{k}(\tau_{k})$. Suppose $\delta\in\left(0,\frac{1}{2}\right)$ satisfies $\delta\max\limits_{1\leq k\leq q}\mu_{k}<\frac{1}{2}$, where $\mu_{k}=\sum_{\theta\in(0,2\pi)}S_{M_{k}}^{-}(e^{\sqrt{-1}\theta})$. Suppose that $\hat{i}(\gamma_{k},1)>0$ for any $k=1,\dots,q$, then there exist infinitely many $(N,m_{1},\dots,m_{q})\in\mathbb{N}^{q+1}$ such that
$I(k,m_{k})=N+\Delta_{k},$
for any $k=1,\dots,q$, where
$I(k,m_{k})=m_{k}\left(\hat{i}(\gamma_{k},1)+S_{M_{k}}^{+}(1)-C(M_{k})\right)+\sum_{\theta\in(0,2\pi)}E\left(\frac{m_{k}\theta}{\pi}\right)S_{M_{k}}^{-}(e^{\sqrt{-1}\theta}),\ \ \Delta_{k}=\sum_{0<\left\{m_{k}\frac{\theta}{\pi}\right\}<\delta}S_{M_{k}}^{-}(e^{\sqrt{-1}\theta}).$
Morover, we have
$\min\left\{\left\{\frac{m_{k}\theta}{\pi}\right\},1-\left\{\frac{m_{k}\theta}{\pi}\right\}\right\}<\delta,$ (2.7)
$m_{k}\frac{\theta}{\pi}\in\mathbb{N},\quad\text{if}\ \ \frac{\theta}{\pi}\in\mathbb{Q},$ (2.8)
where $e^{\sqrt{-1}\theta}\in\sigma(M_{k})$, $\frac{\theta}{\pi}\in(0,2)$ and $\delta$ can be chosen as small as we want.
According to [19], the existence of infinitely many integer tuples $(N,m_{1},\dots,m_{q})\in\mathbb{N}^{q+1}$ in Theorem 2.1 can be transformed into a dynamical system problem on the torus. Let
$h=q+\sum_{k=1}^{q}\mu_{k}$
and
$\nu=\left(\frac{1}{\hat{i}(\gamma_{1,1})},\dots,\frac{1}{\hat{i}(\gamma_{q,1})},\frac{\theta_{1,1}}{\pi\hat{i}(\gamma_{1,1})},\frac{\theta_{1,2}}{\pi\hat{i}(\gamma_{1,1})},\dots,\frac{\theta_{1,\mu_{1}}}{\pi\hat{i}(\gamma_{1,1})},\frac{\theta_{2,1}}{\pi\hat{i}(\gamma_{2,1})},\dots,\frac{\theta_{q,\mu_{q}}}{\pi\hat{i}(\gamma_{q,1})}\right)\in\mathbb{R}^{h},$ (2.9)
where $e^{\sqrt{-1}\theta_{i,j}}\in\sigma(M_{i})$ for $1\leq j\leq\mu_{i}$, then Theorem 2.1 is equivalent to find a vertex
$\chi=(\chi_{1},\dots,\chi_{q},\chi_{1,1},\dots,\chi_{1,\mu_{1}},\dots\chi_{q,1},\dots,\chi_{q,\mu_{q}})$
on the torus $0,1]^{h}$ and infinitely many integers $N\in\mathbb{N}$ such that
$\bigl|\{N\nu\}-\chi\bigr|<\epsilon,$ (2.10)
where $\epsilon\in(0,\min\{\delta,1/3\})$, and $\chi_{k},\chi_{k,j}=0$ or $1$ for $k=1,\dots,q$ and $j=1,\dots,\mu_{1},2,\dots,\mu_{2},\dots,\mu_{q}$. In fact, such a vertex $\chi$ exists, as stated in Theorem 2.2.
Theorem 2.2.
(cf. Theorem 4.2 of [19]) Let the closure of $\{mv\mid m\in\mathbb{N}\}$ on $T^{h}=(\mathbb{R}/\mathbb{Z})^{h}$ be $H$, and let $\pi:\mathbb{R}^{h}\to T^{h}$ be the projection map. Let $V=T_{0}\pi^{-1}H$ denote the tangent space of $\pi^{-1}H$ at the origin of $\mathbb{R}^{h}$, and define
$A(v)=V\setminus\bigcup_{v_{k}\in\mathbb{R}\setminus\mathbb{Q}}\left\{x=(x_{1},\dots,x_{h})\in V\mid x_{k}=0\right\}.$
Define $\psi(x)$ such that $\psi(x)=0$ if $x\geq 0$, and $\psi(x)=1$ if $x<0$. Then for any $a=(a_{1},\dots,a_{h})\in A(v)$, the vector $\chi(a)=\left(\psi(a_{1}),\dots,\psi(a_{h})\right)$ satisfies
$\bigl|\{Nv\}-\chi\bigr|<\epsilon$
for infinitely many $N\in\mathbb{N}$. Moreover, this set $A(v)$ possesses the property: If $v\in\mathbb{R}^{h}\setminus\mathbb{Q}^{h}$, then $\dim V\geq 1,0\notin A(v)\subset V,A(v)=-A(v)$ and $A(v)$ is open in $V$.
Remark 2.3.
Given $M_{0}\in\mathbb{N}$, by the proof of Theorem 4.1 of [19], we may further require $M_{0}|N$ (since the closure of the set $\{\{Nv\}:N\in\mathbb{N},M_{0}|N\}$) is still a closed additive subgroup of $T^{h}$ for some $h\in\mathbb{N}$. Then we can use the step 2 in Theorem 4.1 of of [19] to get $N$), hence in our choice of $(N,m_{1},...,m_{q})$ in Theorem 2.1, we can choose $M_{0}$ good enough such that $N\in\mathbb{N}$ further satisfies
$\frac{N}{M\hat{i}(\gamma_{k},1)}\in\mathbb{Z},\ \ \text{for}\ \forall\ \hat{i}(\gamma_{k},1)\in\mathbb{Q},\ \ k\in\{1,\ldots,q\}.$ (2.11)
Furthermore, from (2.10), we get $\hat{i}(\gamma_{k},1)\in\mathbb{Q}$ implies $\chi_{k}(a)=\psi(a_{k})=0$.
In the following, the Maslov-type index According to [19], for each closed characteristics $(\tau,x)\in\widetilde{\mathcal{J}}(\Sigma,\alpha)$, $\gamma_{x}(\tau)$ is symplectic similar to a matrix with the form $N_{1}(1,1)\diamond M$ with $M\in\mathrm{Sp}(2n-2)$, $\hat{i}(x)>2$ and there is a subset $\mathcal{V}_{\infty}(\Sigma,\alpha)=\{(\tau_{j},x_{j})\,|\,j=1,\cdots,q\}\subset\widetilde{\mathcal{J}}(\Sigma,\alpha)$ such that Theorem 2.1 can be applied to it. Thus, we can obtain a correspondence between closed characteristics and some given integers.
Theorem 2.4.
(cf.[19]) For some $a\in A(v)$, let $(N,m_{1},\cdots,m_{q})\in\mathbb{N}^{q+1}$ be given in Theorem 2.1 and 2.2. Define $\varrho_{n}(\Sigma):\mathcal{H}(2n)\to\mathbb{N}\cup\{+\infty\}$ by
$\varrho_{n}(\Sigma)=\begin{cases}+\infty,&\text{if }^{\#}\mathcal{V}_{\infty}(\Sigma,\alpha)=+\infty,\\ \min\left\{\frac{\hat{i}(x,1)+2S^{+}(x)-\nu(x,1)+n}{2}\mid(\tau,x)\in\mathcal{V}_{\infty}(\Sigma,\alpha)\right\},&\text{if }^{\#}\mathcal{V}_{\infty}(\Sigma,\alpha)<+\infty.\end{cases}$
Then for $s=1,\ldots,\varrho_{n}(\Sigma)$, $x_{j(s)}$ are geometric different, thus there are at least $\varrho_{n}(\Sigma)\geq\lfloor\frac{n}{2}\rfloor+1$ closed characteristics. Moreover, at least $\varrho_{n}(\Sigma)-1\geq[\frac{n}{2}]$ among them have irrational mean index. Also, for each $s=1,\ldots,\varrho_{n}(\Sigma)$, there exists a unique $j(s)\in\{1,\ldots,q\}$ and an injection map $p:\mathbb{N}+K\to\mathcal{V}_{\infty}(\Sigma,\alpha)\times\mathbb{N}$ for some $K\geq 0$, such that $p(N-s+1)=((\tau_{j(s)},x_{j(s)}),2m_{j(s)})$ and
$i(x_{j(s)},2m_{j(s)})\leq 2N-2s+n\leq i(x_{j(s)},2m_{j(s)})+\nu(x_{j(s)},2m_{j(s)})-1.$
Moreover, for any $s_{1},s_{2}\in\{1,\ldots,\varrho_{n}(\Sigma)\}$ with $s_{1}<s_{2}$, we have:
$\hat{i}(x_{j(s_{2})},2m_{j(s_{2})})<\hat{i}(x_{j(s_{1})},2m_{j(s_{1})}).$
and if all the closed characteristics in non-degenerate, then $\varrho_{n}(\Sigma)=n$.
Based on this Theorem, we can further obtain following lemma.
Lemma 2.5.
We can choose an $\hat{a}\in A(v)$ such that $\chi_{j(s)}(\hat{a})=1$ for some $s=1,...,\varrho_{n}(\Sigma)$.
Proof.
Take $a\in A(v)$, from the definition of $A(v)$, one can see that for any i=1,hi=1,...h, if $v_{i}\in\mathbb{R}\setminus\mathbb{Q}$, then $a_{i}\neq 0$. Since $v$ is given by (2.9), then $v_{i}=1/\hat{i}(\gamma_{i,1})$ for i=1,,qi=1,...,q, this combine with Theorem 2.4 tells us that at least $\varrho_{n}(\Sigma)-1\geq[\frac{n}{2}]$ of $v_{j(s)},s=1,...,\varrho_{n}(\Sigma)$ are irrational number. Therefore there exist at least $\varrho_{n}(\Sigma)-1\geq[\frac{n}{2}]$ of $a_{j(s)},s=1,...,\varrho_{n}(\Sigma)$ are non-zero. If $a_{j(s)}<0$ for some $s\in\{1,...,\varrho_{n}(\Sigma)\}$, then $\chi_{j(s)}(a)=1$, take $\hat{a}=a$, such $\hat{a}$ is found. If $a_{j(s)}>0$ for some $s\in\{1,...,\varrho_{n}(\Sigma)\}$, since A(v)=A(v)A(v)=-A(v), we can take $\hat{a}=-a\in A(v)$, then $\hat{a}_{j(s)}<0$, hence $\chi_{j(s)}(\hat{a})=1$, this complete the proof. ∎
3 The elliptic closed characteristics
Based on Theorems 2.1 and Theorem 2.4, Hu-Ou [10] give some inequalities below. These inequalities are useful in our proof of the main Theorems.
Lemma 3.1.
(cf. Thm 2.7 of [10]) For any $s\in\{1,\dots,\varrho_{n}(\Sigma)\}$, we have
$2s\geq n+S_{M_{j(s)}}^{+}(1)+C(M_{j(s)})-2\Delta_{j(s)}-\nu(x_{j(s)},2m_{j(s)})+1,$ (3.1)
$2s\leq n+S_{M_{j(s)}}^{+}(1)+C(M_{j(s)})-2\Delta_{j(s)}.$ (3.2)
Lemma 3.2.
(cf. Cor 2.8 of [10]) For any $s_{1},s_{2}\in\{1,\dots,\varrho_{n}(\Sigma)\}$ with $s_{1}<s_{2}$,
i) If $D_{j(s_{2})}\in\mathbb{Q}$, then $\chi_{j(s_{2})}(a)=0$ and
$N=\left(\left[\frac{N}{MD_{j(s_{2})}}\right]+\chi_{j(s_{2})}(a)\right)MD_{j(s_{2})}$ $<$ $\left(\left[\frac{N}{MD_{j(s_{1})}}\right]+\chi_{j(s_{1})}(a)\right)MD_{j(s_{1})}$
with $D_{j(s_{1})}\in\mathbb{R}\setminus\mathbb{Q}$, $\chi_{j(s_{1})}(a)=1$.
ii) If $\chi_{j(s_{2})}(a)=1$, then $D_{j(s_{2})}\in\mathbb{R}\setminus\mathbb{Q}$ and
$N<\left(\left[\frac{N}{MD_{j(s_{2})}}\right]+\chi_{j(s_{2})}(a)\right)MD_{j(s_{2})}$ $<$ $\left(\left[\frac{N}{MD_{j(s_{1})}}\right]+\chi_{j(s_{1})}(a)\right)MD_{j(s_{1})}$
with $D_{j(s_{1})}\in\mathbb{R}\setminus\mathbb{Q}$, $\chi_{j(s_{1})}(a)=1$.
iii)
$\chi_{j(s_{2})}(a)\leq\chi_{j(s_{1})}(a).$
Based on the $iii)$ of Lemma 3.2, Lemma 2.5, we can easy to obtain below a simple but useful lemma.
Lemma 3.3.
We can choose an $\hat{a}\in A(v)$ such that $\chi_{j(1)}(\hat{a})=1$.
Proof.
[…]
We must mention that the variable $\Delta_{j(s)}$ in Lemma 3.2 is important in the analysis of the ellipticity of the closed characteristics. In the following, we present an non-trivial estimation for variable $\Delta_{j(s)}$.
Theorem 3.4.
For any $s\in\{1,\dots,\varrho_{n}(\Sigma)\}$, we have following estimation
$\Delta_{j(s)}\leq r-\tilde{r}+r_{*}-\tilde{r}_{*}$. (3.3)
Moreover, if $\chi_{j(s)}(a)=0$ and $\hat{i}(x_{j(s)},1)\in\mathbb{R}\setminus\mathbb{Q}$, then we have
$\Delta_{j(s)}\leq r-\tilde{r}-1+r_{*}-\tilde{r}_{*}$. (3.4)
Proof.
From the properties of the splitting number in Section 2.2, for variable $\Delta_{j(s)}$, we have
$\Delta_{j(s)}$ $=$ $\sum_{\{m_{j(s)}\frac{\theta}{\pi}\}\in(0,\delta)}S_{M_{j(s)}}^{-}(e^{\sqrt{-1}\theta})$ (3.5)
$=$ $\sum_{\begin{subarray}{c}l=1\\ \{m_{j(s)}\frac{\theta_{l}}{\pi}\}\in(0,\delta)\end{subarray}}^{r}S_{M_{j(s)}}^{-}(e^{\sqrt{-1}\theta_{l}})+\sum_{\begin{subarray}{c}l=1,\,\theta=\alpha_{l}\,\text{or}\,2\pi-\alpha_{l},\\ \{m_{j(s)}\frac{\theta}{\pi}\}\in(0,\delta)\end{subarray}}^{r_{*}}S_{M_{j(s)}}^{-}(e^{\sqrt{-1}\theta}).$
First, from (2.8) in Theorem 2.1, we have the trivial estimation
$\sum_{\begin{subarray}{c}l=1\\ \{m_{j(s)}\frac{\theta_{l}}{\pi}\}\in(0,\delta)\end{subarray}}^{r}S_{M_{j(s)}}^{-}(e^{\sqrt{-1}\theta_{l}})\leq r-\tilde{r}$. (3.6)
Moreover, one can see that if $\{m_{j(s)}\frac{\alpha_{l}}{\pi}\}\in(0,\delta)$, then $\{m_{j(s)}\frac{2\pi-\alpha_{l}}{\pi}\}=\{1-m_{j(s)}\frac{\alpha_{l}}{\pi}\}\in(1-\delta,1)$. Similar, if $\{m_{j(s)}\frac{2\pi-\alpha_{l}}{\pi}\}\in(0,\delta)$, then $\{m_{j(s)}\frac{\alpha_{l}}{\pi}\}\in(1-\delta,1)$. Take $\delta$ small enough, only one of $\{m_{j(s)}\frac{\alpha_{l}}{\pi}\}$, $\{m_{j(s)}\frac{2\pi-\alpha_{l}}{\pi}\}$ belong to $(0,\delta)$, hence we have estimation
$\sum_{\begin{subarray}{c}l=1,\,\theta=\alpha_{l}\,\text{or}\,2\pi-\alpha_{l},\\ \{m_{j(s)}\frac{\theta}{\pi}\}\in(0,\delta)\end{subarray}}^{r_{*}}S_{M_{j(s)}}^{-}(e^{\sqrt{-1}\theta})\leq r_{*}-\tilde{r}_{*}$. (3.7)
Then inequalities (3.6), (3.7) and formula (3.5) implies (3.3).
Now, assume that $\chi_{j(s)}(a)=0$ and $\hat{i}(x_{j(s)},1)\in\mathbb{R}\setminus\mathbb{Q}$, then (2.4) implies that at least one of $\frac{\theta_{1}}{\pi},\frac{\theta_{2}}{\pi},\ldots,\frac{\theta_{r}}{\pi}$ is irrational number, hence $r-\tilde{r}\geq 1$ and
$\left\{m_{j(s)}D_{j(s)}\right\}$ $=$ $\left\{m_{j(s)}\left(i(x_{j(s)},1)+p_{-}+p_{0}-r+\sum_{l=1}^{r}\frac{\theta_{l}}{\pi}\right)\right\}$
$=$ $\left\{m_{j(s)}\sum_{l=1,\,\frac{\theta_{l}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}}^{r}\frac{\theta_{j}}{\pi}\right\}$
$\leq$ $\sum_{l=1,\,\frac{\theta_{l}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}}^{r}\left\{m_{j(s)}\frac{\theta_{j}}{\pi}\right\},$ (3.9)
where $m_{j(s)}=([\frac{N}{MD_{j(s)}}]+\chi_{j(s)}(a))M=[\frac{N}{MD_{j(s)}}]M$. The second equality comes from (2.8) in Theorem 2.5.
On the other hand,
$\left\{m_{j(s)}D_{j(s)}\right\}$ $=$ $\left\{\left[\frac{N}{MD_{j(s)}}\right]MD_{j(s)}\right\}$ (3.10)
$=$ $\left\{N-\left\{\frac{N}{MD_{j(s)}}\right\}MD_{j(s)}\right\},$
and from (2.9), (2.10) and $\chi_{j(s)}(a)=0$, we get that $\{\frac{N}{MD_{j(s)}}\}=\left|\{\frac{N}{MD_{j(s)}}\}-\chi_{j(s)}(a)\right|<\epsilon$ for any given $\epsilon$ small enough, take $0<\epsilon<\frac{1-r\delta}{MD_{j(s)}}$ with $\delta$ small enough, where $\delta$ is given by (2.7) in Theorem 2.1, then from (3.10), we obtain $\{m_{j(s)}D_{j(s)}\}>r\delta,$ this combines with (3.9), we have
$r\delta<\sum_{l=1,\,\frac{\theta_{l}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}}^{r}\left\{m_{j(2)}\frac{\theta_{j}}{\pi}\right\},$ (3.11)
hence at least one of the elements in $\{\frac{\theta_{j}}{\pi}|\frac{\theta_{j}}{\pi}\in\mathbb{R}\setminus\mathbb{Q},j=1,\ldots,r\}$ satisfies $\{m_{j(s)}\frac{\theta_{j}}{\pi}\}\not\in(0,\delta)$. This combine with the properties of splitting number in Section 2.2, we have
$\sum_{l=1,\{m_{j(s)}\frac{\theta_{l}}{\pi}\}\in(0,\delta)}^{r}S_{M_{j(s)}}^{-}(e^{\sqrt{-1}\theta_{l}})\leq r-\tilde{r}-1,$ (3.12)
Combine with (3.12), (3.5), (3.6) and (3.7), we the following estimation, this completes the proof,
$\Delta_{j(s)}\leq r-\tilde{r}-1+r_{*}-\tilde{r}_{*}$.
Based on Theorem 3.4 and inequalities 3.1 in Lemma 3.2, we have following estimation.
Corollary 3.5.
For any $s\in\{1,\dots,\varrho_{n}(\Sigma)\}$, we have
$2s\geq p_{-}+q_{+}+2r_{*}+2(r_{0}-\tilde{r}_{0})+k+1.$ (3.13)
Moreover, if $\chi_{j(s)}(a)=0$ and $\hat{i}(x_{j(s)},1)\in\mathbb{R}\setminus\mathbb{Q}$, we further have
$2s\geq p_{-}+q_{+}+2r_{*}+2(r_{0}-\tilde{r}_{0})+k+3.$ (3.14)
Proof.
Based on (2.8) in Theorem 2.1, we can rewrite formula (2.3) by
$\nu(\gamma_{j(s)},2m_{j(s)})=p_{-}+2p_{0}+p_{+}+q_{-}+2q_{0}+q_{+}+2(\tilde{r}+\tilde{r}_{*}+\tilde{r}_{0}).$
This combine with formula (2.5), (2.6) and inequality (3.3), (3.4), we obtain (3.13), (3.14). ∎
Using the above inequalities, we obtain the following important result.
Theorem 3.6.
The closed characteristics $x_{j(1)},x_{j(2)}$ has following properties:
(1) $x_{j(1)}$ is an elliptic closed characteristic and
$\gamma_{j(1)}(\tau)$ $\simeq N_{1}(1,1)\diamond I_{2p_{0}}\diamond N_{1}(1,-1)^{\diamond p_{+}}\diamond N_{1}(-1,1)^{\diamond q_{-}}\diamond-I_{2q_{0}}$ (3.15)
$\ \ \diamond R(\theta_{1})\diamond\cdots\diamond R(\theta_{r})\diamond N_{2}(\lambda_{1},v_{1})\diamond\cdots\diamond N_{2}(\lambda_{r_{0}},v_{r_{0}}),$
with $r_{0}=\tilde{r}_{0}$ (i.e.all $N_{2}(\lambda_{1},\nu_{1}),\ldots,N_{2}(\lambda_{r_{0}},\nu_{r_{0}})$ are rational normal form). Moreover, when $\Sigma$ is non-degenerate, then $x_{j(1)}$ is irrationally elliptic.
(2)   $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}\Leftrightarrow\chi_{j(1)}(a)=1$.
(3) If  $\hat{i}(x_{j(2)},1)\in\mathbb{R}\setminus\mathbb{Q}$ and $\chi_{j(2)}(a)=0$, then $x_{j(2)}$ is an elliptic closed characteristic and it admits a symplectic normal form of similar type (3.15). When $\Sigma$ is non-degenerate, then $x_{j(2)}$ is irrationally elliptic.
Proof.
For $(1)$, from the fact $p_{-}\geq 1$ and inequality (3.13) in Corollary 3.5, when $s=1$, we have
$2\geq p_{-}+q_{+}+2r_{*}+2(r_{0}-\tilde{r}_{0})+k+1\geq 2+q_{+}+2r_{*}+2(r_{0}-\tilde{r}_{0})+k.$
hence $p_{-}=1,q_{+}=r_{*}=k=0$ and $r_{0}=\tilde{r}_{0}$, from the symplectic decomposition (2.2), we get (3.15). Moreover, when the compact convex surface $\Sigma$ is non-degenerate, we further have $p_{0}=p_{+}=q_{-}=q_{0}=\tilde{r}=\tilde{r}_{0}=0$, hence $x_{j(1)}$ is irrationally elliptic.
For $(2)$, if $\hat{i}(x_{j(1)},1)\in\mathbb{Q}$ then from $i)$ in Lemma 3.2, we must have $\chi_{j(1)}(a)=0$, hence $\chi_{j(1)}(a)=1$ implies $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$. Moreover, if  $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$ and $\chi_{j(1)}(a)=0$, using inequality (3.14) in Corollary 3.5, when $s=1$, we have
$2\geq p_{-}+q_{+}+2r_{*}+2(r_{0}-\tilde{r}_{0})+k+3\geq 4+q_{+}+2r_{*}+2(r_{0}-\tilde{r}_{0})+k,$
which implies $2\geq 4$, contradiction. Therefore, $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$ must implies $\chi_{j(1)}(a)=1$.
For $(3)$, if $\chi_{j(2)}(a)=0$ and $\hat{i}(x_{j(2)},1)\in\mathbb{R}\setminus\mathbb{Q}$, using inequality (3.14) in Corollary 3.5, when $s=2$, we have
$4\geq p_{-}+q_{+}+2r_{*}+2(r_{0}-\tilde{r}_{0})+k+3\geq 4+q_{+}+2r_{*}+2(r_{0}-\tilde{r}_{0})+k.$
hence $p_{-}=1,q_{+}=r_{*}=k=0$ and $r_{0}=\tilde{r}_{0}$, from the symplectic decomposition (2.2), then $x_{j(2)}$ have a symplectic normal form of type (3.15). Similar, when the compact convex surface $\Sigma$ is non-degenerate, we further have $p_{0}=p_{+}=q_{-}=q_{0}=\tilde{r}=\tilde{r}_{0}=0$, hence $x_{j(2)}$ is irrationally elliptic. ∎
Based on the above Theorem 3.6, we can easy to prove our main Theorem 1.2.
Proof of Theorem 1.2..
For any fix $a\in A(v)$ and $s\in\{1,\ldots,\varrho_{n}(\Sigma)\}$, consider the inject map $p(N-s+1)=([(\tau_{j(s)},x_{j(s)})],2m_{j(s)})$ given by Theorem 2.4. Now, from Lemma 3.3, we can choose an $\hat{a}\in A(v)$ such that $\chi_{j(1)}(\hat{a})=1$. Assume that the compact convex surface $\Sigma$ is non-degenerate, then $(1)$ in Theorem 3.6 implies $x_{j(1)}$ is irrationally elliptic. Since $\chi_{j(1)}(\hat{a})=1$, from $(2)$ in Theorem 3.6, we know $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$. Now, choose $\tilde{a}=-\hat{a}\in A(v)$, using Theorem 2.4 again, we still have $(\tilde{N},\tilde{m}_{1},\ldots,\tilde{m}_{q})$ , $\tilde{j}(s)$ and inject map $p(\tilde{N}-s+1)=([(\tau_{\tilde{j}(s)},x_{\tilde{j}(s)})],2\tilde{m}_{\tilde{j}(s)}),s\in\{1,\ldots,\varrho_{n}(\Sigma)\}$. Similar, $(1)$ in Theorem 3.6 implies $x_{\tilde{j}(1)}$ is irrationally elliptic. If $\tilde{j}(1)\neq j(1)$, then $x_{j(1)},x_{\tilde{j}(1)}$ are two different irrationally elliptic closed characteristics. If $\tilde{j}(1)=j(1)$, from the definition of $\chi(a)$, we know that $\chi_{j(1)}(\hat{a})=1$ implies $\chi_{j(1)}(-\hat{a})=0$, hence $\chi_{\tilde{j}(1)}(\tilde{a})=\chi_{j(1)}(-\hat{a})=0$, but $\hat{i}(x_{\tilde{j}(1)},1)=\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$, this contradict with $(2)$ in Theorem 3.6, hence this case will not happen, we always have $\tilde{j}(1)\neq j(1)$ and we get two irrationally elliptic closed characteristics. This completes the proof. ∎
Moreover, under Assumption A, by a similar proof of Theorem 1.2, we can easily obtain Theorem 1.5.
Proof of Theorem 1.5..
From Lemma 3.3, we can choose an $\hat{a}\in A(v)$ such that $\chi_{j(1)}(\hat{a})=1$, Assumption A tells us that $x_{j(1)}$ is irrational elliptic , then similar the proof of Theorem 1.2, take $a=-\hat{a}$, we have inject map $p(\tilde{N}-s+1)=([(\tau_{\tilde{j}(s)},x_{\tilde{j}(s)})],2\tilde{m}_{\tilde{j}(s)}),s\in\{1,\ldots,\varrho_{n}(\Sigma)\}$, then $x_{\tilde{j}(1)}$ must distinct from $x_{j(1)}$, use Assumption A again, $x_{\tilde{j}(1)}$ is another irrational elliptic closed characteristics. ∎
In the following, we will prove main Theorem 1.6, that is the existence of three elliptic closed characteristics on any $C^{2}$ compact convex hypersurfaces $\Sigma\subset\mathbb{R}^{6}$. First, We need the following useful result, which is a corollary of Theorem A as proved by Çineli-Ginzburg-Gürel-Mazzucchellid [2].
Corollary 3.7.
(cf. Theorem A of [2]) For any $C^{2}$ compact convex hypersurface $\Sigma\subset\mathbb{R}^{2n}$, if $\Sigma$ has a hyperbolic closed characteristics, then $\Sigma$ has infinitely many geometrically distinct closed characteristics.
Remark 3.8.
The original Theorem A in [2] is more general, the authors consider the closed Reeb orbits on the (2n1)(2n-1)-dimensional standard contact sphere, under very mild dynamical convexity type assumptions, the presence of one hyperbolic closed orbit implies the existence of infinitely many simple closed Reeb orbits. It well known that the Hamiltonian flow on the hypersurface $\Sigma$ is equivalent to the Reeb flow on a standard contact sphere. Moreover, when $\Sigma$ is convex, it satisfies the dynamical convexity type assumptions in [2], therefore the above result is just a simple corollary of Theorem A in [2].
[…]
$S_{M}^{+}(1)=p_{-}=1,\ \ C(M)=r+2r_{*}.$
Moreover, Lemma 3.2 shows that for any $s\in\{1,\dots,\varrho_{n}(\Sigma)\}$, we have $\nu(x_{j(s)},2m_{j(s)})=1$, hence the inequalities in Lemma 3.2 become equalities
$2s$ $=$ $n+S_{M_{j(s)}}^{+}(1)+C(M_{j(s)})-2\Delta_{j(s)}$ (3.16)
$=$ $4+r+2r_{*}-2\Delta_{j(s)}.$
Based on this equalities and Corollary 3.7, we can obtain three elliptic closed characteristics in the following.
Proof of Theorem 1.6..
First, Theorem 3.6 implies $x_{j(1)}$ is irrationally elliptic. For $s=2$, formula (3.16) becomes $0=r+2r_{*}-2\Delta_{j(2)}$, from Corollary 3.7, we know that $x_{j(2)}$ can not be hyperbolic. Since $\Sigma$ is non-degenerate, if $x_{j(2)}$ is not an elliptic closed characteristics, from the symplectic decomposition (2.2), the only possible case is
$\gamma_{j(2)}(\tau)$ $\simeq N_{1}(1,1)\diamond R(\theta_{1})\diamond M_{1},$
then $r=1$, this contradict with $0=r+2r_{*}-2\Delta_{j(2)}$, hence $x_{j(2)}$ must be elliptic. For $s=3$, formula (3.16) becomes $2=r+2r_{*}-2\Delta_{j(2)}\leq r+2r_{*}$, hence $r=2$ or $r_{*}=1$, this implies $x_{j(3)}$ is elliptic. Therefore we have proved the existence of three elliptic closed characteristics.
Furthermore, Theorem 2.4 tells us that at least $\varrho_{3}(\Sigma)-1=2$ among $x_{j(1)},x_{j(2)}$ and $x_{j(3)}$ have irrational mean index, hence we obtain that one of $\hat{i}(x_{j(2)})$ and $\hat{i}(x_{j(3)})$ must be irrational. Without loss of generality, we assume that $\hat{i}(x_{j(2)})$ is irrational, since we have proved that $x_{j(2)}$ is elliptic, then from the mean index formula (2.4), the only possible case is
$\gamma_{j(2)}(\tau)$ $\simeq N_{1}(1,1)\diamond R(\theta_{1})\diamond R(\theta_{2}).$
Since $\Sigma$ is non-degenerate, $x_{j(2)}$ must be irrational elliptic. In summary, we obtain that $x_{j(1)}$ is irrational elliptic and either $x_{j(2)}$ or $x_{j(3)}$ must be irrational elliptic. This completes the proof. ∎
Acknowledgments. The authors would like to express the gratitude to Xijun Hu for helpful discussions, and particularly for his constructive guidance of the research direction. This work is partially supported by the Taishan Scholars Climbing Program of Shandong($\sharp$ TSPD20240802). Y.Ou is partially supported by NSFC ($\sharp$12371192), the Young Taishan Scholars Program of Shandong Province ($\sharp$ tsqn202312055), and the Qilu Young Scholar Program of Shandong University.

### 16 Mar 2026 v2

@@
 Even in $\mathbb{R}^{6}$, it is unknown whether there exist three elliptic closed characteristics.
-In this paper, we first prove that for any non-degenerate $\Sigma\in\mathbb{R}^{2n}$ with finitely many closed characteristics, at least two of them are irrationally elliptic.
+In this paper, we first prove that for any $\Sigma\subset\mathbb{R}^{2n}$ with finitely many closed characteristics, there exist at least two elliptic closed characteristics, which possess a nice symplectic normal form.
+In particular, as a simple corollary, they are irrational elliptic when $\Sigma$ is non-degenerate.
 Moreover, for any non-degenerate $\Sigma\subset\mathbb{R}^{6}$ with finitely many closed characteristics, we obtain at least three elliptic characteristics, of which at least two are irrationally elliptic.
@@
 This result provide theoretical support for further research on this conjecture.
-AMS Subject Classification: 58E05, 37J45, 34C25
+AMS Subject Classification: 58E05, 53D12, 34C25
 Key Words: Compact convex hypersurfaces, Closed characteristics, Stability, Maslov-type index, Hamiltonian systems
@@
 The eigenvalues of $\gamma_{x}(\tau)$ are called Floquet multipliers, and denote the elliptic height $e(\gamma_{x}(\tau))$ of $\gamma_{x}(\tau)$ to be the total algebraic multiplicity of all eigenvalues of $\gamma_{x}(\tau)$ on the unit circle $\mathbb{U}=\{z\in\mathbb{C}\,|\,|z|=1\}$.
-As usual, a closed characteristic $(\tau,x)$ is non-degenerate, if $1\in\sigma(x)$ is of algebraic multiplicity $2$; a hypersurface $\Sigma$ is called non-degenerate, if all the closed characteristics on the given hypersurface together with all of their iterations are non-degenerate as periodic solutions of the corresponding Hamiltonian system.
-A closed characteristic $(\tau,x)$ is called elliptic, if $e(\gamma_{x}(\tau))=2n$; hyperbolic if $1$ is a double Floquet multiplier of it and $e(\gamma_{x}(\tau))=2$; irrationally elliptic if in an appropriate coordinate $\gamma_{x}(\tau)$ is the $\diamond$-product of one $N_{1}(1,1)=\left(\begin{array}[]{cc}1&1\\ 0&1\\ \end{array}\right)$ and n1n-1 rotation matrices $\left(\begin{array}[]{cc}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\\ \end{array}\right)$ with rotation angles $\theta_{i}$ being irrational multiples of $\pi$ for $1\leq i\leq n-1$.
+As usual, a closed characteristic $(\tau,x)$ is non-degenerate, if $1\in\sigma(x)$ is of algebraic multiplicity 22; a hypersurface $\Sigma$ is called non-degenerate, if all the closed characteristics on the given hypersurface together with all of their iterations are non-degenerate as periodic solutions of the corresponding Hamiltonian system.
+A closed characteristic $(\tau,x)$ is called elliptic, if $e(\gamma_{x}(\tau))=2n$; hyperbolic if 11 is a double Floquet multiplier of it and $e(\gamma_{x}(\tau))=2$; irrationally elliptic if in an appropriate coordinate $\gamma_{x}(\tau)$ is the $\diamond$-product of one $N_{1}(1,1)=\left(\begin{array}[]{cc}1&1\\ 0&1\\ \end{array}\right)$ and n1n-1 rotation matrices $\left(\begin{array}[]{cc}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\\ \end{array}\right)$ with rotation angles $\theta_{i}$ being irrational multiples of $\pi$ for $1\leq i\leq n-1$.
 Hence, the stability problem refers to estimating the number of elliptic closed characteristics that may exist for system (1.3).
@@
 For results on more general compact star-shaped hypersurfaces, we refer the reader to [12], [9], [13], [11], etc.
-Remark 1.1.
+Remark $1.$
 In general, when $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, it hard to estimate the upper bound of $^{\#}\widetilde{\mathcal{J}}(\Sigma)$. This relate to the following conjecture.
@@
 Theorem 1.2.
-For any non-degenerate surfaces $\Sigma\in\mathcal{H}(2n)$ with $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then at least two closed characteristics are irrational elliptic.
+For any surfaces $\Sigma\in\mathcal{H}(2n)$ with $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then at least two closed characteristics are elliptic whose momodormy matrix possess a nice symplectic normal form
+$\gamma(\tau)$ $\simeq N_{1}(1,1)\diamond I_{2p_{0}}\diamond N_{1}(1,-1)^{\diamond p_{+}}\diamond N_{1}(-1,1)^{\diamond q_{-}}\diamond-I_{2q_{0}}$
+$\ \ \diamond R(\theta_{1})\diamond\cdots\diamond R(\theta_{r})\diamond N_{2}(\lambda_{1},v_{1})\diamond\cdots\diamond N_{2}(\lambda_{r_{0}},v_{r_{0}}),$
+with all $N_{2}(\lambda_{1},\nu_{1}),\ldots,N_{2}(\lambda_{r_{0}},\nu_{r_{0}})$ are rational normal form. Moreover, when $\Sigma$ is non-degenerate, then they are irrationally elliptic.
 Remark 1.3.
-By combining the methods of [1], [11] and this paper, the result of Theorem 1.2 can perhaps be generalized to dynamically convex Reeb flow on non-degenerate star-shaped surfaces.
+By combining the methods of [1], [11] and this paper, the result of Theorem 1.2 can perhaps be generalized to dynamically convex Reeb flow on star-shaped surfaces.
 Remark 1.4.
-How to remove the non-degeneracy condition in Theorem 1.2 is an interesting question, but for now, we are unable to do so.
+How to obtain the irrationally elliptic without the non-degeneracy condition in Theorem 1.2 is an interesting question, but for now, we are unable to do so.
 In fact, it is still unknown whether the following Assumption A holds, except for the case $n=2$.
@@
 Note that the $\diamond$-product of any two symplectic matrices is symplectic.
-For any two paths $\gamma_{j}\in\mathcal{P}_{\tau}(2n_{j})$ with $j=0$ and $1$, let $\gamma_{0}\diamond\gamma_{1}(t)=\gamma_{0}(t)\diamond\gamma_{1}(t)$ for all $t\in[0,\tau]$.
+For any two paths $\gamma_{j}\in\mathcal{P}_{\tau}(2n_{j})$ with $j=0$ and 11, let $\gamma_{0}\diamond\gamma_{1}(t)=\gamma_{0}(t)\diamond\gamma_{1}(t)$ for all $t\in[0,\tau]$.
 For $\gamma(\tau)\in\mathrm{Sp}(2n)_{\omega}^{*}$, the $\omega$-index of $\gamma$ can be defined by
@@
 […]
-Here we call $N_{2}(\omega,b)$ nontrivial if $(b_{2}-b_{3})\sin\theta<0$ and trivial if $(b_{2}-b_{3})\sin\theta>0$. For any $M_{i}\in\emph{Sp}(2n_{i})$ with $i=0$ and $1$, there holds
+Here we call $N_{2}(\omega,b)$ nontrivial if $(b_{2}-b_{3})\sin\theta<0$ and trivial if $(b_{2}-b_{3})\sin\theta>0$. For any $M_{i}\in\emph{Sp}(2n_{i})$ with $i=0$ and 11, there holds
 $S^{\pm}_{M_{0}\diamond M_{1}}(\omega)=S^{\pm}_{M_{0}}(\omega)+S^{\pm}_{M_{1}}(\omega),\ \ \text{and}\ \ S^{\pm}_{M}(\omega)=S^{\mp}_{M}(\bar{\omega}),\ \ \forall\ \ \omega\in\mathbb{U}.$
@@
 $\bigl|\{N\nu\}-\chi\bigr|<\epsilon,$ (2.10)
-where $\epsilon\in(0,\min\{\delta,1/3\})$, and $\chi_{k},\chi_{k,j}=0$ or $1$ for $k=1,\dots,q$ and $j=1,\dots,\mu_{1},2,\dots,\mu_{2},\dots,\mu_{q}$.
-In fact, such a vertex $\chi$ exists, as stated in Theorem 2.2.
-Theorem 2.2.
+where $\epsilon\in(0,\min\{\delta,1/3\})$, and $\chi_{k},\chi_{k,j}=0$ or 11 for $k=1,\dots,q$ and $j=1,\dots,\mu_{1},2,\dots,\mu_{2},\dots,\mu_{q}$.
+In fact, such a vertex $\chi$ exists, as stated in Theorem $2.$
+Theorem $2.$
 (cf.
@@
 Given $M_{0}\in\mathbb{N}$, by the proof of Theorem 4.1 of [19], we may further require $M_{0}|N$ (since the closure of the set $\{\{Nv\}:N\in\mathbb{N},M_{0}|N\}$) is still a closed additive subgroup of $T^{h}$ for some $h\in\mathbb{N}$.
-Then we can use the step 2 in Theorem 4.1 of of [19] to get $N$), hence in our choice of $(N,m_{1},...,m_{q})$ in Theorem 2.1, we can choose $M_{0}$ good enough such that $N\in\mathbb{N}$ further satisfies
+Then we can use the step 2 in Theorem 4.1 of of [19] to get NN), hence in our choice of $(N,m_{1},...,m_{q})$ in Theorem 2.1, we can choose $M_{0}$ good enough such that $N\in\mathbb{N}$ further satisfies
 $\frac{N}{M\hat{i}(\gamma_{k},1)}\in\mathbb{Z},\ \ \text{for}\ \forall\ \hat{i}(\gamma_{k},1)\in\mathbb{Q},\ \ k\in\{1,\ldots,q\}.$ (2.11)
@@
 Theorem 2.4.
-(cf.[19]) For some $a\in A(v)$, let $(N,m_{1},\cdots,m_{q})\in\mathbb{N}^{q+1}$ be given in Theorem 2.1 and 2.2. Define $\varrho_{n}(\Sigma):\mathcal{H}(2n)\to\mathbb{N}\cup\{+\infty\}$ by
+(cf.[19]) For some $a\in A(v)$, let $(N,m_{1},\cdots,m_{q})\in\mathbb{N}^{q+1}$ be given in Theorem 2.1 and $2.$ Define $\varrho_{n}(\Sigma):\mathcal{H}(2n)\to\mathbb{N}\cup\{+\infty\}$ by
 $\varrho_{n}(\Sigma)=\begin{cases}+\infty,&\text{if }^{\#}\mathcal{V}_{\infty}(\Sigma,\alpha)=+\infty,\\ \min\left\{\frac{\hat{i}(x,1)+2S^{+}(x)-\nu(x,1)+n}{2}\mid(\tau,x)\in\mathcal{V}_{\infty}(\Sigma,\alpha)\right\},&\text{if }^{\#}\mathcal{V}_{\infty}(\Sigma,\alpha)<+\infty.\end{cases}$
@@
 Based on the $iii)$ of Lemma 3.2, Lemma 2.5, we can easy to obtain below a simple but useful lemma.
-Lemma 3.3.
+Lemma $3.$
 We can choose an $\hat{a}\in A(v)$ such that $\chi_{j(1)}(\hat{a})=1$.
@@
 and from (2.9), (2.10) and $\chi_{j(s)}(a)=0$, we get that $\{\frac{N}{MD_{j(s)}}\}=\left|\{\frac{N}{MD_{j(s)}}\}-\chi_{j(s)}(a)\right|<\epsilon$ for any given $\epsilon$ small enough, take $0<\epsilon<\frac{1-r\delta}{MD_{j(s)}}$ with $\delta$ small enough, where $\delta$ is given by (2.7) in Theorem 2.1, then from (3.10), we obtain $\{m_{j(s)}D_{j(s)}\}>r\delta,$ this combines with (3.9), we have
-$r\delta<\sum_{l=1,\,\frac{\theta_{l}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}}^{r}\left\{m_{j(2)}\frac{\theta_{j}}{\pi}\right\},$ (3.11)
+$r\delta<\sum_{l=1,\,\frac{\theta_{l}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}}^{r}\left\{m_{j(s)}\frac{\theta_{j}}{\pi}\right\},$ (3.11)
 hence at least one of the elements in $\{\frac{\theta_{j}}{\pi}|\frac{\theta_{j}}{\pi}\in\mathbb{R}\setminus\mathbb{Q},j=1,\ldots,r\}$ satisfies $\{m_{j(s)}\frac{\theta_{j}}{\pi}\}\not\in(0,\delta)$.
@@
 Now, from Lemma 3.3, we can choose an $\hat{a}\in A(v)$ such that $\chi_{j(1)}(\hat{a})=1$.
-Assume that the compact convex surface $\Sigma$ is non-degenerate, then $(1)$ in Theorem 3.6 implies $x_{j(1)}$ is irrationally elliptic.
-Since $\chi_{j(1)}(\hat{a})=1$, from $(2)$ in Theorem 3.6, we know $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$.
+Then Theorem 3.6-$(1)$ implies $x_{j(1)}$ is an elliptic closed characteristic and
+$\gamma_{j(1)}(\tau)$ $\simeq N_{1}(1,1)\diamond I_{2p_{0}}\diamond N_{1}(1,-1)^{\diamond p_{+}}\diamond N_{1}(-1,1)^{\diamond q_{-}}\diamond-I_{2q_{0}}$
+$\ \ \diamond R(\theta_{1})\diamond\cdots\diamond R(\theta_{r})\diamond N_{2}(\lambda_{1},v_{1})\diamond\cdots\diamond N_{2}(\lambda_{r_{0}},v_{r_{0}}),$
+with all $N_{2}(\lambda_{1},\nu_{1}),\ldots,N_{2}(\lambda_{r_{0}},\nu_{r_{0}})$ are rational normal form.
+Moreover, it is easy to see that $x_{j(1)}$ is irrationally elliptic when $\Sigma$ is non-degenerate.
+Since $\chi_{j(1)}(\hat{a})=1$, from Theorem 3.6-$(2)$, we know $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$.
 Now, choose $\tilde{a}=-\hat{a}\in A(v)$, using Theorem 2.4 again, we still have $(\tilde{N},\tilde{m}_{1},\ldots,\tilde{m}_{q})$ , $\tilde{j}(s)$ and inject map $p(\tilde{N}-s+1)=([(\tau_{\tilde{j}(s)},x_{\tilde{j}(s)})],2\tilde{m}_{\tilde{j}(s)}),s\in\{1,\ldots,\varrho_{n}(\Sigma)\}$.
-Similar, $(1)$ in Theorem 3.6 implies $x_{\tilde{j}(1)}$ is irrationally elliptic.
-If $\tilde{j}(1)\neq j(1)$, then $x_{j(1)},x_{\tilde{j}(1)}$ are two different irrationally elliptic closed characteristics.
-If $\tilde{j}(1)=j(1)$, from the definition of $\chi(a)$, we know that $\chi_{j(1)}(\hat{a})=1$ implies $\chi_{j(1)}(-\hat{a})=0$, hence $\chi_{\tilde{j}(1)}(\tilde{a})=\chi_{j(1)}(-\hat{a})=0$, but $\hat{i}(x_{\tilde{j}(1)},1)=\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$, this contradict with $(2)$ in Theorem 3.6, hence this case will not happen, we always have $\tilde{j}(1)\neq j(1)$ and we get two irrationally elliptic closed characteristics.
+Similarly, by applying Theorem 3.6-$(1)$ again, we obtain $x_{\tilde{j}(1)}$ has the same properties as $x_{j(1)}$.
+If $\tilde{j}(1)=j(1)$, from the definition of $\chi(a)$, we know that $\chi_{j(1)}(\hat{a})=1$ implies $\chi_{j(1)}(-\hat{a})=0$, hence $\chi_{\tilde{j}(1)}(\tilde{a})=\chi_{j(1)}(-\hat{a})=0$, but $\hat{i}(x_{\tilde{j}(1)},1)=\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$, this contradict with Theorem 3.6-$(2)$, hence this case will not happen, we always have $\tilde{j}(1)\neq j(1)$, then $x_{j(1)},x_{\tilde{j}(1)}$ are two different closed characteristics that possess the desired properties as required in Theorem 1.2.
 This completes the proof.
@@
 Without loss of generality, we assume that $\hat{i}(x_{j(2)})$ is irrational, since we have proved that $x_{j(2)}$ is elliptic, then from the mean index formula (2.4), the only possible case is
-$\gamma_{j(2)}(\tau)$ $\simeq N_{1}(1,1)\diamond R(\theta_{1})\diamond R(\theta_{2}).$
-Since $\Sigma$ is non-degenerate, $x_{j(2)}$ must be irrational elliptic.
+$\gamma_{j(2)}(\tau)$ $\simeq N_{1}(1,1)\diamond R(\theta_{1})\diamond\hat{M}$,
+where $\hat{M}=R(\theta_{2}),N(\pm 1,a)$ with $a=\pm 1,0$.
+Since $\Sigma$ is non-degenerate, $\hat{M}$ must be $R(\theta_{2})$ with $\frac{\theta_{2}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}$, then $x_{j(2)}$ must be irrational elliptic.
 In summary, we obtain that $x_{j(1)}$ is irrational elliptic and either $x_{j(2)}$ or $x_{j(3)}$ must be irrational elliptic.